OpenStudy (anonymous):

Determine x if (xsin^2 30°cos^2 60°)/(4cos^2 45°) =( 3sin ^2 45°+2cos^2 30°) /(sin^2 90°-4cos^2 45°)

OpenStudy (anonymous):

Do you know where to start?

OpenStudy (anonymous):

Sorry?

OpenStudy (anonymous):

The first thing to do is simplify all the sins & cos. sin 30 = 1/2 cos 60 = 1/2 cos 45 = Sqrt(1/2) sin 45 = Sqrt(1/2) Cos 30 = Sqrt(3) / 2 Sin 90 = 1

OpenStudy (anonymous):

so sin^2 30 = 1/4 cos^2 60 = 1/4 cos^2 45 = 1/2 sin^2 45 = 1/2 cos^2 30 = 3/4 sin^2 90 = 1 Plug those in, and simplify, solve for x

OpenStudy (anonymous):

the sin, cos, and tan of 0, 30, 45, 60 & 90 will be something you just need to memorize.

OpenStudy (anonymous):

Thanks!!! Will memorise them ASAP. I understand that the ratios are the basics for trigonometry. Thanks a lot!!